ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 230, 2019

Aichinger, E. Bounding the free spectrum of nilpotent algebras of prime power order

919

Alonso-Gutiérrez, D. and J. Bernués The square negative correlation property on ℓ_{p}^{n}- balls

895

Banakh, T. and E. Jabłońska Null-finite sets in topological groups and their applications

361

Bernués, J. See D. Alonso-Gutiérrez Blok, R. J., C. G. Hoffman and S. V. Shpectorov Classification

of Curtis–Tits and Phan amalgams with 3-spherical diagram

97

Bocea, M. and M. Mihăilescu On the existence and uniqueness of exponentially harmonic maps and some related problems

795

Brieussel, J. and T. Zheng Shalom’s property H_{FD} and extensions by ℤ of locally finite groups

45

Callegaro, F. and M. Salvetti Homology of the family of hyperelliptic curves

653

Chaika, J. and D. Constantine Quantitative shrinking target properties for rotations and interval exchanges

275

Chernikov, A. and N. Hempel Mekler’s construction and generalized stability

745

Clark, L. O., D. Martín Barquero, C. Martín González and M. Siles Molina Using the Steinberg algebra model to determine the center of any Leavitt path algebra

23

Conant, G. Stability and sparsity in sets of natural numbers  471

Constantine, D. See J. Chaika Cordes, M. and D. Hume Relatively hyperbolic groups with fixed

peripherals

443

Cúth, M., O. F. K. Kalenda and P. Kaplický Finitely additive measures and complementability of Lipschitz-free spaces

409

Domínguez Benavides, T., M. A. Japón and J. Villada Bedoya Linearly and directionally bounded weak-star closed sets and the AFPP

509

Eskin, A. and C. Matheus Semisimplicity of the Lyapunov spectrum for irreducible cocycles

973

Finis, T. and E. Lapid On the analytic properties of intertwining operators II: Local degree bounds and limit multiplicities 771 Glasby, S. P., C. E. Praeger and B. Xia ‘Norman involutions’

and tensor products of unipotent Jordan blocks

153

Golubev, K. and O. Parzanchevski Spectrum and combinatorics of two-dimensional Ramanujan complexes

583

Grandjean, V. Re-parameterizing and reducing families of normal operators

715

Hachtman, S. Determinacy and monotone inductive definitions  71

Hempel, N. See A. Chernikov Hoffman, C. G. See R. J. Blok Hume, D. See M. Cordes Jabłońska, E. See T. Banakh Jaoui, R. Differential fields and geodesic flows II: Geodesic flows of

pseudo-Riemannian algebraic varieties

527

Japón, M. A. See T. Domínguez Benavides Kalai, G. and L. J. Schulman Quasi-random multilinear polyno-

mials

195

Kalenda, O. F. K. See M. Cúth Kaplický, P. See M. Cúth

Kastis, E. The norm closed triple semigroup algebra  855
Kramer, K. and R. Miller The Hilbert’s-Tenth-Problem Operator  693

Kwiatkowska, A. and M. Malicki Automorphism groups of countable structures and groups of measurable functions

335

Lapid, E. See T. Finis Leung, D. H., L. Li, T. Oikhberg and M. A. Tursi Separable

universal Banach lattices

141

Li, L. See D. H. Leung Lüders, C. M. and C. Reiher The Ramsey–Turán problem for

cliques

613

Málek, M. and S. Roth Constant slope models and perturbation  213

Malicki, M. See A. Kwiatkowska Mangoubi, D. and A. Weller Weiser Harmonic functions vanish-

ing on a cone

563

Martín Barquero, D. See L. O. Clark Martín González, C. See L. O. Clark Matheus, C. See A. Eskin McGoff, K. and R. Pavlov Factor maps and embeddings for ran-

dom ℤ^{d} shifts of finite type

239

Mermelstein, O. Calculating the closed ordinal Ramsey number

R^{cl}(ω · 2, 3) 387

Metsch, K. An Erdős–Ko–Rado theorem for finite buildings of

type F_{4} 813

Mihăilescu, M. See M. Bocea Miller, R. See K. Kramer Oikhberg, T. See D. H. Leung Parzanchevski, O. See K. Golubev Pavlov, R. See K. McGoff Praeger, C. E. See S. P. Glasby Reiher, C. See C. M. Lüders Roth, S. See M. Málek Salvetti, M. See F. Callegaro Schulman, L. J. See G. Kalai

Shmerkin, P. On the Hausdorff dimension of pinned distance sets  949

Shpectorov, S. V. See R. J. Blok Shumyatsky, P., P. Zalesskii and T. Zapata Profinite groups in

which centralizers are abelian

831

Siles Molina, M. See L. O. Clark Todorov, I. G. and L. Turowska Transference and preservation

of uniqueness

1

Tsukamoto, M. Mean dimension of full shifts  183

Turowska, L. See I. G. Todorov Tursi, M. A. See D. H. Leung Villada Bedoya, J. See T. Domínguez Benavides Weller Weiser, A. See D. Mangoubi Xia, B. See S. P. Glasby Zalesskii, P. See P. Shumyatsky Zapata, T. See P. Shumyatsky Zheng, T. See J. Brieussel